Searches related to Consider the full-wave bridge rectifier circuit. The input signal is 120 V (rms) at 60 Hz. The load resistance is RL = 250 . The peak output voltage is to be 9 V and the ripple voltage is to be no more than 5 percent. Determine the required turns ratio and the load current.
1
Expert's answer
2020-12-10T09:40:25-0500

Determine the peak value "v_{s(max)}"

"v_{s(max)} = v_{0(max)} + 2V_\u03b3"

"V_\u03b3" is the diode cut-in voltage

"v_{0(max)}" is the desired peak value of the output voltage

"V_γ = 0.7 \;V \\

v_{0(max)} = 9 \;V \\

v_{s(max)} = 9 + 2 \times 0.7 = 10.4 \;V"

Determine the root mean square of the output sinusoid signal:

"v_{s,rms} = \frac{v_{s(max)}}{\sqrt{2}} \\

= \frac{10.4}{\sqrt{2}} = 7.35 \;V"

Determine the required turn ratio:

"k = \frac{N_1}{N_2} = \frac{v_{I,rms}}{v_{s,rms}}"

"N_1" is the number of winding on the primary side of the transformer

"N_2" is the number of winding on the secondary side of the transformer

"v_{I,rms}" is the root mean square of the input voltage

"v_{s,rms}" is the root mean square of the sinusoid output signal

"v_{I,rms} = 120 \;V(rms) \\

k = \frac{120}{7.35} = 16.3"

The required turns ratio is 16.3